How much 35% copper alloy should be melted with 37 kg of 75% copper alloy to produce an alloy which is 50% copper. If necessary round your answer to the nearest tenth.
Answer by checkley77(12844)

.75*37+.35X=.50(37+X)
27.75+.35X=18.5+.50X
.35X-.50X=18.5-27.75
-.15X=-9.25
X=-9.25/-.15
X=61.7 KG OF 35% COPPER IS USED.
PROOF:
.75*37+.35*61.7=.50(37+61.7)
27.8+21.6=.50*98.7
49.4=49.49.4